# Multitask Learning Can Improve Worst-Group Outcomes

### 05 Dec 2023 v1

@@ 3.3 Training on auxiliary data only
$\mathcal{L_{\mathrm{recon}}}\left(W^{\mathrm{aux}},a\right)=\frac{1}{2M}\sum_{(\tilde{x}_{i},x_{i})\in\mathbf{T}_{\mathrm{aux}}}\|x_{i}-\big{(}W^{\mathrm{aux}}\big{)}^{T}f(\tilde{x})\|^{2}$
Figure 3: The ratio $\log(\frac{\mathbf{a}_{\mathrm{spur}}}{\mathbf{a}_{\mathrm{core}}})$ for 2 (extreme) choices of $|\mathbf{a}|_{1}$ across 4 hyperparameter settings (learning rate $\times$ batch size). $\log(\frac{\mathbf{a}_{\mathrm{spur}}}{\mathbf{a}_{\mathrm{core}}})$ < 0 means that the spurious features are assigned lower weight compared to the core ones.
Figure 3: The ratio $\log(\frac{\mathbf{a}_{\mathrm{spur}}}{\mathbf{a}_{\mathrm{core}}})$ for 2 (extreme) choices of $|\mathbf{a}|_{1}$ across 4 hyperparameter settings (learning rate $\times$ batch size). $\log(\frac{\mathbf{a}_{\mathrm{spur}}}{\mathbf{a}_{\mathrm{core}}})$ < 0 means that the spurious features are assigned lower weight compared to the core ones.
@@
Using the synthetic data instantiation in Figure 1, we apply noise from $\mathcal{N}\left(\mathbf{0}\leavevmode\nobreak\ ,\leavevmode\nobreak\ I_{2}\right)$ – i.e $\sigma^{2}_{\text{noise}}=1$ – on each of the 1,000 training points to get the training data for the auxiliary task ***Note that whilst we could generate more points for the auxiliary task, we would like to mimic the setting where we refrain from introducing external data (data beyond end task training data), since methods we will compare against like JTT do not introduce external data.. We fit the model solely to the auxiliary task by running batched stochastic gradient descent on $\mathcal{L_{\mathrm{recon}}}$. We use learning rates in the set $\{10^{-2},10^{-3}\}$ and batch sizes in the set $\{64,256\}$.
We consider two cases, when the intermediate layer $\mathbf{a}$ has low versus high capacity as reflected in $\ell_{1}$-norm. Low $\ell_{1}$-norm – $\|\mathbf{a}\|_{1}=|\mathbf{a}_{\text{spur}}|+|\mathbf{a}_{\text{core}}|=0.1$ means restricted capacity since this constraint (along with $\|W^{\mathrm{aux}}\|_{F}=1$) results in models that cannot fit the training data perfectly. High $\ell_{1}$-norm ($\|\mathbf{a}\|_{1}=10$) means that the model is expressive enough to perfectly fit the training data. Figure 4 provides some insight about the learned intermediate layer in either case. When the model has enough capacity, there is no competition between the core and spurious features. This means solutions where the spurious feature is weighted more than the core feature are feasible as long as the core feature weight is enough to reconstruct the the noised core features well.
However, under restricted capacity where the learned weight of the core and spurious features are in direct competition, the model has to put more weight on the core features in order to achieve a lower auxiliary loss (as motivated in Section 2). This can be seen from Figure 4. Thus, for the auxiliary task to be effective at forcing a model to use core features over the spurious ones, the capacity of the model has to be reasonably restricted.
@@ 3.4 Multitasking with regularization
We implement the multitask model illustrated in Figure 4. We use the same set of hyper-parameters as used in Section 3.3 and perform joint stochastic gradient descent on both $\mathbf{T}_{\text{end}}$ and $\mathbf{T}_{\text{aux}}$.
When $\tau$ is chosen to be small enough, model capacity is restricted and the model is forced to chiefly rely on the core features in order to do well on both the end and auxiliary tasks. Figure 5 evinces this. When the norm of the shared layer is high $\|\mathbf{a}\|_{1}=10$, the end task can still predominantly rely on the spurious features leading to poor worst group error. On the other hand, when model capacity is reasonably restricted by setting $\|\mathbf{a}\|_{1}=0.1$, we see from Figure 5 (left) that we can achieve improved worst group accuracy. Thus, in this simplified setting, we are able to effectively leverage the reconstruction auxiliary task by applying sufficient regularization to ensure improved worst group outcomes.
@@
Figure 5: Depicted are the learned half-spaces for the multitask model under $\tau=\{0.1,10\}$ and $\alpha=10$. Restricting the capacity of the shared feature space is critical for multitasking to be effective for improving worst group error. Examples visualized are 1,000 balanced test examples sampled from Equation 1
@@ 4 Details For Natural Data Experiments
We have made a case for using regularized multitask learning as an intervention for combating poor worst-group performance through empirical explorations in a simplified, synthetic setting. In this section, we go over the experimental details for our investigations of tasks of more practical interest.
@@ 4.1 Datasets
We conduct experiments across three datasets. To relieve the burden of compute, we introduce a fourth dataset which is a smaller, sub-sampled version of one of the original datasets for ablations.
Waterbirds:
This image classification dataset was introduced by Sagawa et al. (2020a). The task is to distinguish between species of land and water birds. It consists of bird images sourced from the CUB dataset Wah et al. (2011) and superimposed on land or water backgrounds from the Places dataset Zhou et al. (2018). The label (type of bird) is spuriously correlated with background, resulting in 4 groups. Since this is a small dataset (4,795 train examples), we also use it for ablations.
MultiNLI:
This is a natural language inference dataset. The task is to classify whether the second sentence is entailed by, contradicts, or is neutral with respect to the first sentence (Williams et al., 2018). Following Sagawa et al. (2020a), we utilize the presence of negation words as a spurious attribute, leading to the creation of a total of 6 groups.
Civilcomments:
The Civilcomments dataset is a toxicity classification dataset that contains comments from online forums Borkan et al. (2019); Koh et al. (2021). Along with the toxicity label, each text is annotated with additional overlapping sub-group labels of 8 demographic identities: male, female, LGBTQ, Christian, Muslim, other religions, Black, and White. As per Koh et al. (2021) and Sagawa et al. (2020a), we defines 16 overlapping groups by taking the Cartesian product of the binary toxicity label and each of the above 8 demographic identities.
Civilcomments-small:
As Civilcomments is a large dataset of about 448,000 datapoints, we create a sub-group stratified subset of 5% for conducting ablations and other detailed experiments. Our subset contains 13,770, 2,039, and 4,866 datapoints in our train, validation, and test split, respectively.
@@ 4.2 Multitask Model and Training Details
We follow the parameter sharing paradigm (Ruder, 2017; Sener & Koltun, 2018) where both $\mathbf{T}_{\mathrm{end}}$ and $\mathbf{T}_{\mathrm{aux}}$ share the same model body, parameterized by $\theta_{\mathrm{base}}$. We instantiate task-specific heads, parameterized by $\theta_{\mathrm{end}}$ and $\theta_{\mathrm{aux}}$, respectively. We introduce $\ell_{1}$ regularization to the last layer activations immediately before the per-task prediction heads. Specifically, let $h^{\mathrm{end}}$, $h^{\mathrm{aux}}\in\mathbb{R}^{d}$ be the output representations generated by the base model, which are fed into their respective task-specific heads. Our final multitask learning objective is expressed as follows
@@
For training, we vary the fine-tuning learning rate within the ranges of $\{10^{-3},10^{-4}\}$ for Waterbirds, and $\{10^{-4},10^{-5}\}$ for the text datasets. We experiment with batch sizes in the set $\{4,8,16,32\}$ for all the datasets. We use the same batch sizes for $\mathbf{T}_{\mathrm{end}}$ and $\mathbf{T}_{\mathrm{aux}}$. We train for $50$ epoch for the NLP datasets and $200$ epochs for Waterbirds, with an early stopping patience of $10$, as per the check-pointing scheme explained in section 4.2. We cross-validate $\alpha_{\mathrm{aux}}$ and $\alpha_{\mathrm{reg}}$ from $\{e^{-1},1,e^{1}\}$. We use the Adam optimizer for NLP datasets with decoupled weight decay regularization of $10^{-2}$ (Loshchilov & Hutter, 2017). Consistent with the recent studies on ViT (Dosovitskiy et al., 2020; Steiner et al., 2022), we use SGD with momentum of $0.9$ (Sutskever et al., 2013) to fine-tune Waterbirds. We run each hyperparameter configuration across $5$ seeds and report the averaged results. We report the ERM, JTT, and groupDRO results for Civilcomments and MultiNLI from Idrissi et al. (2022) as the authors conducted extensive hyperparameter tuning across all these methods. However, since Idrissi et al. (2022) report results on Waterbirds using a ResNet-50 model (He et al., 2016) and our experiments employ ViT, we re-run all baselines using ViT with a consistent set of hyperparameters, as mentioned above.
@@
We assess all methods and datasets using worst-group accuracy and employ two model selection strategies:
Val-GP:
This strategy requires group annotations in the validation data during training. Here, we checkpoint the model based on the maximum worst-group accuracy on the validation data.
No-GP:
This strategy requires no access to any group annotations during training. We checkpoint models based on the average validation accuracy.
@@
Empirical Risk Minimization (ERM):
This is the standard approach of minimizing the average loss over all the training data. No group information is used during training except when the Val-GP strategy is used for model selection.
Just Train Twice (JTT):
It presents a two step approach for worst group generalization Liu et al. (2021). JTT first trains a standard ERM model for $T$ epochs to identify misclassified datapoints. Then, a second model is trained on a reweighted dataset constructed by up weighting the mis-classified examples by $\alpha_{up}$. It does not use group information during training except for the Val-GP strategy.
Bit-rate Constrained DRO (BR-DRO):
Traditionally, in the two-player formulation of DRO, the adversary can use complex re-weighting functions, resulting in overly pessimistic solutions. In contrast, BR-DRO Setlur et al. (2023) constrains the adversary’s complexity based on information theory under a data-independent prior. While BR-DRO offers weaker robustness without performance guarantees for arbitrary re-weighting, it is less pessimistic and suitable for simpler distribution shifts, characterized by re-weighting function contained in a simpler complexity class. BR-DRO does not use group information during training except for the Val-GP setting.
Group-DRO:
Group distributionally robust optimization minimizes the maximum loss across all the sub-groups Sagawa et al. (2020a). This optimization method incorporates group annotations during training. Similar to prior works Liu et al. (2021); Idrissi et al. (2022); Setlur et al. (2023), we treat it as an oracle, as this is the only method that uses group annotations.
@@ 5 Results And Discussion
In this section, we provide empirical evidence demonstrating the effectiveness of our regularized multitask learning approach in mitigating worst-group performance while maintaining average performance across different scenarios. Unless explicitly stated otherwise, the auxiliary tasks employed for multitasking exclusively rely on end-task data only.
@@ 5.1 Multitasking is Competitive with Bespoke DRO Methods
We first compare our approach with previously proposed methods for tackling worst-group accuracy. Table 3 details the performance of various methods across the tasks of interest for the Val-GP setting. As expected, groupDRO yields the highest worst-group accuracy as it directly optimizes for it. Our MTL approach outperforms JTT and BR-DRO on two datasets (MNLI and Waterbirds) whilst performing comparatively with BR-DRO on the CivilComments dataset. Given the competitive results in Table 3, we argue that our regularized MTL formulation is an attractive option over JTT and BR-DRO. Multitasking already features prominently in many ML code bases. Thus, introducing our simple regularization modification to existing MTL implementations represents smaller technical overhead compared to introducing JTT or BR-DRO to target worst-group error. Also, as we will see in the Section 5.2 below, regularized MTL is a single approach which is capable of improving both worst-group and average accuracy.
@@ 5.2 Multitasking Improves both Average and worst-group Performance Even in the Absence Group Annotations
Though previous works typically assume that practitioners have access to the group annotations on the validation set (Liu et al., 2021; Kirichenko et al., 2022), we are interested in settings where no such annotations are available. This covers many tasks of practical interest since, in some cases, it may be prohibitively cost-intensive (financially and in terms of human labor) to acquire group annotations even for the smaller validation set (Paranjape et al., 2023). Consequently, we present a comparative performance analysis in Figure 6, encompassing settings with and without access to group annotations. With respect to worst-group accuracy, our regularized MTL approach outperforms JTT and achieves $\approx 2\%$ lift over ERM when group annotations are absent, a trend consistent across both Waterbirds and Civilcomments-small datasets. Whilst this lift of $\approx 2\%$ remains when validation group annotations are introduced, the benefit from group-labeled data is more pronounced ($\approx 5\%-15\%$). This boost can be worthwhile to practitioners who have the resources to obtain some group annotations. Moreover, it becomes evident from Figure 6 that our method not only yields superior worst-group performance but also improves in average performance.
@@ 5.3 Are both Regularization and Multitasking Jointly necessary?
Our approach proposes multitasking the end task with an auxiliary task while regularizing the final layer joint embedding space. We conduct an ablation to verify if both ingredients are necessary to improve average and worst-group performance. Our results are captured in Table 4. When assessing the worst-group accuracy, we find that regularizing the final embedding space during ERM can, at times, result in worse performance compared to training via standard ERM ($66.3_{1.6}$ vs $67.4_{2.1}$ on CivilComments-small with validation group labels). On the other hand, multitasking without regularization can fail to improve over ERM, as evinced by the lack of improvement on Waterbirds. The regularized MTL approach is the only setting that consistently improves on both datasets with and without validation group annotations. Consistent with these findings, we observe that joint L1 regularization and multitask learning yield the highest average accuracy in most cases.

### 28 Feb 2024 v2

@@ 3.3 Training on auxiliary data only
$\mathcal{L_{\mathrm{recon}}}\left(W^{\mathrm{aux}},a\right)=\frac{1}{2M}\sum_{(\tilde{x}_{i},x_{i})\in\mathbf{T}_{\mathrm{aux}}}\|x_{i}-\big{(}W^{\mathrm{aux}}\big{)}^{T}f(\tilde{x})\|^{2}$ (7)
Figure 3: The ratio $\log(\frac{\mathbf{a}_{\mathrm{spur}}}{\mathbf{a}_{\mathrm{core}}})$ for 2 (extreme) choices of $|\mathbf{a}|_{1}$ across 4 hyperparameter settings (learning rate $\times$ batch size).
Figure 3: The ratio $\log(\frac{\mathbf{a}_{\mathrm{spur}}}{\mathbf{a}_{\mathrm{core}}})$ for 2 (extreme) choices of $|\mathbf{a}|_{1}$ across 4 hyperparameter settings (learning rate $\times$ batch size).
@@
Using the synthetic data instantiation in Figure 1, we apply noise from $\mathcal{N}\left(\mathbf{0}\leavevmode\nobreak\ ,\leavevmode\nobreak\ I_{2}\right)$, i.e., $\sigma^{2}_{\text{noise}}=1$ on each of the $1000$ training points to get the training data for the auxiliary task ***Note that while we could generate more points for the auxiliary task, we would like to mimic the setting where we refrain from introducing external data (data beyond end task training data) since methods like JTT and BR-DRO do not utilize them.. We fit the model solely to the auxiliary task by running batched stochastic gradient descent on $\mathcal{L_{\mathrm{recon}}}$. We use learning rates in the set $\{10^{-2},10^{-3}\}$ and batch sizes in the set $\{64,256\}$.
We consider two cases when the intermediate layer $\mathbf{a}$ has low versus high capacity, as reflected in $\ell_{1}$-norm. Low $\ell_{1}$-norm – $\|\mathbf{a}\|_{1}=|\mathbf{a}_{\text{spur}}|+|\mathbf{a}_{\text{core}}|=0.1$ means restricted capacity since this constraint (along with $\|W^{\mathrm{aux}}\|_{F}=1$) results in models that cannot fit the training data perfectly. High $\ell_{1}$-norm ($\|\mathbf{a}\|_{1}=10$) means that the model is expressive enough to perfectly fit the training data. Figure 4 provides insight into the learned intermediate layer in either case. When the model has enough capacity, there is no competition between the core and spurious features. This means solutions where the spurious feature is weighted more than the core feature are feasible as long as the core feature weight is enough to reconstruct the the noised core features well.
However, under restricted capacity where the learned weight of the core and spurious features are in direct competition, the model has to put more weight on the core features in order to achieve a lower auxiliary loss (as motivated in Section 2). This can be seen in Figure 4. Thus, for the auxiliary task to be effective at forcing a model to use core features over spurious ones, the model’s capacity must be reasonably restricted.
@@ 3.4 Multitasking with regularization
We implement the multitask model illustrated in Figure 4. We use the same set of hyper-parameters as used in Section 3.3 and perform joint stochastic gradient descent on both $\mathbf{T}_{\text{end}}$ and $\mathbf{T}_{\text{aux}}$. When $\tau$ is chosen to be small enough, model capacity is restricted, and the model is forced to rely chiefly on the core features to do well on both the end and auxiliary tasks. Figure 5 evinces this. When the norm of the shared layer is high $\|\mathbf{a}\|_{1}=10$, the end task can still predominantly rely on the spurious features leading to poor worst group error. On the other hand, when model capacity is reasonably restricted by setting $\|\mathbf{a}\|_{1}=0.1$, we see from Figure 5 (left) that we can achieve improved worst group accuracy. Thus, we can effectively leverage the reconstruction auxiliary task in this simplified setting by applying sufficient regularization to ensure improved worst-group outcomes.
@@
Figure 5: Depicted are the learned half-spaces for the multitask model under $\tau=\{0.1,10\}$ and $\alpha=10$. Restricting the capacity of the shared feature space is critical for multitasking to be effective for improving worst group error. Examples visualized are $1000$ balanced test examples sampled from Equation 1
@@ 4 Details For Natural Data Experiments
We have made a case for using regularized multitask learning to combat poor worst-group performance through empirical explorations in a simplified, synthetic setting. In this section, we review the experimental details for our investigations of tasks of more practical interest.
@@ 4.1 Datasets
We conduct experiments across three datasets. To relieve the burden of compute, we introduce a fourth dataset, a smaller, sub-sampled version of one of the original datasets for ablations.
  1. 1.
    Waterbirds: This image classification dataset was introduced by Sagawa et al. (2020a). The task is to distinguish between species of land and water birds. It consists of bird images sourced from the CUB dataset Wah et al. (2011) and superimposed on land or water backgrounds from the Places dataset Zhou et al. (2018). The label (type of bird) is spuriously correlated with the background, resulting in 4 groups. Since this is a small dataset ($4795$ train examples), we also use it for ablations.
  2. 2.
    MultiNLI: This is a natural language inference dataset. The task is to classify whether the second sentence is entailed by, contradicts, or is neutral with respect to the first sentence (Williams et al., 2018). Following Sagawa et al. (2020a), we utilize the presence of negation words as a spurious attribute, leading to the creation of a total of $6$ groups.
  3. 3.
    Civilcomments: The Civilcomments dataset is a toxicity classification dataset that contains comments from online forums Borkan et al. (2019); Koh et al. (2021). Along with the toxicity label, each text is annotated with additional overlapping sub-group labels of 8 demographic identities: male, female, LGBTQ, Christian, Muslim, other religions, Black, and White. As per Koh et al. (2021) and Sagawa et al. (2020a), we defines $16$ overlapping groups by taking the Cartesian product of the binary toxicity label and each of the above $8$ demographic identities.
  4. 4.
    Civilcomments-small: As Civilcomments is a large dataset of about $448000$ datapoints, we create a sub-group stratified subset of $5\%$ for conducting ablations and other detailed experiments. Our subset contains $13770$, $2039$, and $4866$ datapoints in our train, validation, and test split, respectively.
@@ 4.2 Multitask Model and Training Details
We follow the parameter sharing paradigm (Ruder, 2017; Sener & Koltun, 2018) where both $\mathbf{T}_{\mathrm{end}}$ and $\mathbf{T}_{\mathrm{aux}}$ share the same model body, parameterized by $\theta_{\mathrm{base}}$. We instantiate task-specific heads, parameterized by $\theta_{\mathrm{end}}$ and $\theta_{\mathrm{aux}}$, respectively. We introduce $\ell_{1}$ regularization to the last layer activations immediately before the per-task prediction heads ***In contrast to synthetic data experiments, where the norm constraint is applied to pre-prediction layer weights, in this section we directly apply the norm constraint to the features, indirectly constraining model weights. In the synthetic experiment, core and spurious features have a one-to-one mapping to model weights, enabling direct regularization of the features, but that is no longer possible in more complex models.. Specifically, let $h^{\mathrm{end}}$, $h^{\mathrm{aux}}\in\mathbb{R}^{d}$ be the output representations generated by the base model, which are fed into their respective task-specific heads. Our final multitask learning objective is expressed as follows:
@@
For training, we vary the fine-tuning learning rate within the ranges of $\{10^{-3},10^{-4}\}$ for Waterbirds, and $\{10^{-4},10^{-5}\}$ for the text datasets. We experiment with batch sizes in the set $\{4,8,16,32\}$. We use the same batch sizes for $\mathbf{T}_{\mathrm{end}}$ and $\mathbf{T}_{\mathrm{aux}}$. We train for $50$ epoch for the NLP datasets and $200$ epochs for Waterbirds, with an early stopping patience of $10$, as per the check-pointing scheme explained in section 4.2. We use the Adam optimizer for NLP datasets with decoupled weight decay regularization of $10^{-2}$ (Loshchilov & Hutter, 2017). Consistent with the recent studies on ViT (Dosovitskiy et al., 2020; Steiner et al., 2022), we use SGD with a momentum of $0.9$ (Sutskever et al., 2013) to fine-tune Waterbirds. We run each hyperparameter configuration across $5$ seeds and report the averaged results. We report the ERM, JTT, and groupDRO results for Civilcomments and MultiNLI from Idrissi et al. (2022) as the authors conducted extensive hyperparameter tuning across all these methods. However, since Idrissi et al. (2022) report results on Waterbirds using a ResNet-50 model (He et al., 2016) and our experiments employ ViT, we re-run all baselines using ViT with a consistent set of hyperparameters, as mentioned above.
@@
We assess all methods and datasets using two model selection strategies:
  1. 1.
    Val-GP: This strategy requires group annotations in the validation data during training. Here, we select the model based on the maximum worst-group accuracy on the validation data.
  2. 2.
    No-GP: This strategy requires no access to any group annotations during training. We select the model based on the average validation accuracy.
@@
  1. 1.
    Empirical Risk Minimization (ERM): This is the standard approach of minimizing the average loss over all the training data. No group information is used during training except when the Val-GP strategy is used for model selection.
  2. 2.
    Just Train Twice (JTT): It presents a two step approach for worst group generalization Liu et al. (2021). JTT first trains a standard ERM model for $T$ epochs to identify misclassified datapoints. Then, a second model is trained on a reweighted dataset constructed by upweighting the misclassified examples by $\alpha_{up}$. It does not use group information during training except for the Val-GP strategy.
  3. 3.
    Bit-rate Constrained DRO (BR-DRO): Traditionally, in the two-player formulation of DRO, the adversary can use complex reweighting functions, resulting in overly pessimistic solutions. In contrast, BR-DRO Setlur et al. (2023) constrains the adversary’s complexity based on information theory under a data-independent prior. While BR-DRO offers weaker robustness without performance guarantees for arbitrary reweighting, it is less pessimistic and suitable for simpler distribution shifts, characterized by a reweighting function contained in a simpler complexity class. BR-DRO does not use group information during training except for the Val-GP setting.
  4. 4.
    Group-DRO: Group distributionally robust optimization minimizes the maximum loss across all the sub-groups Sagawa et al. (2020a). This optimization method incorporates group annotations during training. Similar to prior works Liu et al. (2021); Idrissi et al. (2022); Setlur et al. (2023), we treat it as an oracle, as this is the only method that uses group annotations.
@@ 5 Results And Discussion
In this section, we provide empirical evidence demonstrating the effectiveness of our regularized MTL approach in mitigating worst-group error while maintaining average performance across different scenarios.
@@ 5.1 Multitasking is Competitive with Bespoke DRO Methods
We first compare our approach with previously proposed methods for tackling worst-group accuracy. Table 3 details the performance of various methods across the tasks of interest for the Val-GP setting. As expected, groupDRO yields the highest worst-group accuracy as it directly optimizes for it. Our MTL approach outperforms JTT and BR-DRO on two datasets (MNLI and Waterbirds) while performing comparatively with BR-DRO on the CivilComments dataset. Given the competitive results in Table 3, we argue that our regularized MTL formulation is an attractive option over JTT and BR-DRO. Multitasking already features prominently in many ML code bases. Thus, introducing our simple regularization modification to existing MTL implementations represents a smaller technical overhead compared to introducing JTT or BR-DRO to target worst-group error. Also, as we will see in Section 5.2 below, regularized MTL is a single approach capable of improving both worst-group and average accuracy.
@@ 5.2 Multitasking Improves both Average and worst-group Performance Even in the Absence Group Annotations
Though previous works typically assume that practitioners have access to the group annotations on the validation set (Liu et al., 2021; Kirichenko et al., 2022), we are interested in settings where no such annotations are available. This covers many tasks of practical interest since, in some cases, it may be prohibitively cost-intensive (financially and in terms of human labor) to acquire group annotations even for the smaller validation set (Paranjape et al., 2023). Consequently, we present a comparative performance analysis in Figure 6, encompassing settings with and without access to group annotations. With respect to worst-group accuracy, our regularized MTL approach outperforms JTT and achieves $\approx 2\%$ lift over ERM when group annotations are absent, a trend consistent across both Waterbirds and Civilcomments-small datasets. While this lift of $\approx 2\%$ remains when validation group annotations are introduced, the benefit from group-labeled data is more pronounced ($\approx 5\%-15\%$). This boost can be worthwhile to practitioners who have the resources to obtain some group annotations. Moreover, it becomes evident from Figure 6 that our method not only yields superior worst-group performance but also improves in average performance.
@@ 5.3 Are both Regularization and Multitasking Jointly necessary?
In this section, we conduct an ablation to verify if both multitask learning and regularizing the final layer joint embedding space are necessary to improve average and worst-group performance. Our results are captured in Table 4. When assessing the worst-group accuracy, we find that regularizing the final embedding space during ERM can, at times, result in worse performance compared to training via standard ERM ($66.3_{1.6}$ vs $67.4_{2.1}$ on CivilComments-small with validation group labels). On the other hand, multitasking without regularization can fail to improve over ERM, as evinced by the lack of improvement on Waterbirds. The regularized MTL approach is the only setting consistently improving on both datasets with and without validation group annotations. In line with these findings, we observe that joint L1 regularization and multitask learning setup yields the highest average accuracy in most cases.